OpenStudy (epicteatime):

How many grams of sodium metal are needed to react completely with 25.8 liters of chlorine gas at 293 Kelvin and 1.30 atmospheres?

OpenStudy (epicteatime):

can people tell me how 2 get moles?

OpenStudy (epicteatime):

@iPwnBunnies @ParthKohli

OpenStudy (epicteatime):

@aaronq

OpenStudy (epicteatime):

@AccessDenied

OpenStudy (epicteatime):

@jigglypuff314

OpenStudy (epicteatime):

can you tell me how to get moles?

OpenStudy (aaronq):

use the ideal gas equation, PV=nRT, to get moles, n.

OpenStudy (epicteatime):

so?

OpenStudy (epicteatime):

how do i do it?

OpenStudy (epicteatime):

n=PV*RT?

OpenStudy (epicteatime):

or divide...;(

OpenStudy (aaronq):

nope, divide n=PV/RT

OpenStudy (epicteatime):

oh

OpenStudy (epicteatime):

P=pressure, V=volume, and R=constant, and T=temp. right?

OpenStudy (aaronq):

correct, T needs to be in Kelvin, btw.

OpenStudy (epicteatime):

oh?

OpenStudy (epicteatime):

n=1.3 x 25.8/0.0821 x 293?

OpenStudy (epicteatime):

is it this?

OpenStudy (aaronq):

yeah (i just noticed the temp is already given in Kelvin lol)

OpenStudy (epicteatime):

lol

OpenStudy (epicteatime):

so is that correct?

OpenStudy (aaronq):

Yeah, now to find the grams of sodium, you need to find the moles of sodium first, then convert.

OpenStudy (epicteatime):

1.294g/mol.

OpenStudy (epicteatime):

whoops, 1.394

OpenStudy (epicteatime):

right?

OpenStudy (aaronq):

yeah, but the units aren't "g/mol", it's just moles

OpenStudy (epicteatime):

oh

OpenStudy (epicteatime):

so i just do the thing, except i add n instead of =n right?

OpenStudy (aaronq):

im not sure i follow

OpenStudy (epicteatime):

?

OpenStudy (epicteatime):

what do i do after i find mole?

OpenStudy (aaronq):

you need to use the molar ratio between them, so write and balance the chemical equation for the process

OpenStudy (epicteatime):

i know the balanced equation 2 Na (s) + Cl2 (g) → 2 NaCl (s)

OpenStudy (epicteatime):

i think its that, it was already given

OpenStudy (aaronq):

yeah that's it. so now, make a ratio of the moles with the coefficients: \(\dfrac{n_{Na}}{2}=\dfrac{n_{Cl_2}}{1}\) plug in moles of what you know, then solve

OpenStudy (epicteatime):

ok...

OpenStudy (epicteatime):

ummm, 1.394Na/2=1.394Cl^2/1?

OpenStudy (epicteatime):

which is 0.697=1.394

OpenStudy (aaronq):

you only know the moles of \(Cl_2\), right? (you found this above) this \(n_{Cl_2}\) is just a symbol for that. so it's \(\dfrac{n_{Na}}{2}=\dfrac{1.394}{1}\rightarrow n_{Na}=\dfrac{1.394*2}{1}\)

OpenStudy (epicteatime):

yeah

OpenStudy (epicteatime):

oh

OpenStudy (epicteatime):

2.788Na

OpenStudy (aaronq):

yeah so that's moles, 2.788 moles of sodium metal now convert the moles to grams, use \(n=\dfrac{m}{M}\)

OpenStudy (epicteatime):

ok

OpenStudy (epicteatime):

1.394=whats m/M?

OpenStudy (aaronq):

m = mass in grams (what you're looking for) M= molar mass (of sodium)

OpenStudy (epicteatime):

ok

OpenStudy (epicteatime):

1.394=m/2.788

OpenStudy (epicteatime):

m=2.788 x 1.394?

OpenStudy (epicteatime):

m=3.89

OpenStudy (aaronq):

the molar mass of sodium is 22.9898 g/mol

OpenStudy (aaronq):

you can find this on a periodic table

OpenStudy (epicteatime):

oh

OpenStudy (epicteatime):

the atomic number?

OpenStudy (epicteatime):

well, the number below

OpenStudy (aaronq):

nope, the atomic number is only the number of protons

OpenStudy (aaronq):

yeah, it's called atomic mass (or molar mass because it's the mass of 1 mole of that substance)

OpenStudy (epicteatime):

ok, so now

OpenStudy (epicteatime):

1.395=22.989/2.788

OpenStudy (aaronq):

hm you're confusing the variables. n=moles = 2.788 moles m=mass (in grams) = x M= molar mass = 22.9898 g/mol the formula is: \(n=\dfrac{m}{M}\)

OpenStudy (epicteatime):

ohhhh

OpenStudy (epicteatime):

2.788=x/22.9898

OpenStudy (epicteatime):

so, x=2.788 x 22.9898?

OpenStudy (aaronq):

yes

OpenStudy (epicteatime):

x=64.1

OpenStudy (aaronq):

\(\checkmark \)

OpenStudy (aaronq):

dont forget the units

OpenStudy (epicteatime):

yay! 64.1g/mol

OpenStudy (aaronq):

nope, just grams. remember this is the mass, not the molar mass

OpenStudy (epicteatime):

kk